• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name R. Subedi et al

    Log entry time 21:57:35 on November 03, 2009

    Entry number 297308


    keyword=DAQ and misc Hall work today

    PMT replacement:

    Bogdan and I replaced all s1 left PMTs in the right HRS. We plan to
    replace 3 more in the same s1 tomorrow. Those three will be top 2 and the
    bottom one in the right side. Also one PMT in S2 in rihgt hrs (probably
    the right side bottom one) will be replaced. In the right hrs gas-cer, we
    will replace a HV connector in the base (it is loose, not sticking tight
    in the base) of right side PMT (2nd from the bottom); currently a HV can
    not be applied to this PMT.

    RHRS detectors:

    Both gas-cer and leadglass detectors are in detector stack and both (left
    and right hrs) stacks are pushed in so as to be ready for hot checkout.
    We have connected gas-cer and preshower cables (shower cable connection
    will be done in a while). Essentially, from detector point of view we
    will be ready in an hour.

    Todo list: Need to finish RHRS trigger setup (tomorrow), currently only
    T1 and T8 are there. Left HRS is already fully functional.